Ticket Number: HEA-NOC/20020708-01 Ticket Status: OPEN
Ticket Type: Scheduled Ticket Source: HEA-NOC
Ticket Scope: Line Site/Line: HEAnet <==> EPA
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: 20020708 11:52:22 UTC Problem Start: 20020708 1400 UTC+1
Ticket Closed: Problem End:
Upgrade of EPA connection to HEAnet.
Will require rehoming of EPA router to new termination point
All traffic to and from EPA.
Mailed will be queued offsite.
Time to Fix:
Outage should be around 1 minute, but could be longer